# Lotus: Diffusion-based Visual Foundation Model for High-quality Dense Prediction

### v4

@@ Abstract
Leveraging the visual priors of pre-trained text-to-image diffusion models offers a promising solution to enhance zero-shot generalization in dense prediction tasks. However, existing methods often uncritically use the original diffusion formulation, which may not be optimal due to the fundamental differences between dense prediction and image generation. In this paper, we provide a systemic analysis of the diffusion formulation for the dense prediction, focusing on both quality and efficiency. And we find that the original parameterization type for image generation, which learns to predict noise, is harmful for dense prediction; the multi-step noising/denoising diffusion process is also unnecessary and challenging to optimize. Based on these insights, we introduce Lotus, a diffusion-based visual foundation model with a simple yet effective adaptation protocol for dense prediction. Specifically, Lotus is trained to directly predict annotations instead of noise, thereby avoiding harmful variance. We also reformulate the diffusion process into a single-step procedure, simplifying optimization and significantly boosting inference speed. Additionally, we introduce a novel tuning strategy called detail preserver, which achieves more accurate and fine-grained predictions. Without scaling up the training data or model capacity, Lotus achieves SoTA performance in zero-shot depth and normal estimation across various datasets. It also enhances efficiency, being significantly faster than most existing diffusion-based methods. Lotus’ superior quality and efficiency also enable a wide range of practical applications, such as joint estimation, single/multi-view 3D reconstruction, etc. Project page: lotus3d.github.io.
Figure 1: We present Lotus, a diffusion-based visual foundation model for dense geometry prediction. With minimal training data, Lotus achieves SoTA performance in two key geometry perception tasks, i.e., zero-shot depth and normal estimation. “Avg. Rank” indicates the average ranking across all metrics, where lower values are better. Bar length represents the amount of training data used.
@@ 1 Introduction
To validate Lotus, we conduct extensive experiments on two primary geometric dense prediction tasks: zero-shot monocular depth and normal estimation. The results demonstrate that Lotus achieves SoTA performance on these tasks across a wide range of evaluation datasets. Compared to traditional discriminative methods, Lotus delivers remarkable results with only 59K training samples by effectively leveraging the powerful diffusion priors. Among generative approaches, Lotus also outperforms previous methods in both accuracy and efficiency, being significantly faster than methods like Marigold (Ke et al., 2024) (Fig. 3). Beyond these improvements, Lotus seamlessly supports various applications, such as joint estimation, single/multi-view 3D reconstruction, etc.
@@
  • Based on our insights, we introduce Lotus, a diffusion-based visual foundation model for dense prediction with simple yet effective fine-tuning protocol. Lotus achieves SoTA performance on both zero-shot monocular depth and surface normal estimation. It also enables a wide range of applications.
@@ 2.2 Generative Models for Dense Perception
Currently, a notable trend involves adopting pre-trained generative models, particularly diffusion models, into dense prediction tasks. Marigold (Ke et al., 2024) and GeoWizard (Fu et al., 2024) directly apply the standard diffusion formulation and the pre-trained parameters, without addressing the inherent differences between image generation and dense prediction, leading to constrained performance. Their efficiency is also severely limited by standard iterative denoising processes and ensemble inferences. In this paper, we propose a novel diffusion formulation tailored to the characteristics of dense prediction. Aiming to fully leveraging the pre-trained diffusion’s powerful visual priors, Lotus enables more accurate and efficient predictions, finally achieving SoTA performance.
More recent works, GenPercept (Xu et al., 2024) and StableNormal (Ye et al., 2024), also adopted single-step diffusion. However, GenPercept (Xu et al., 2024) first removes noise input for deterministic characteristic based on DMP (Lee et al., 2024), and then adopts one-step strategy to avoid surface texture interference. It lacks systematic analysis of the diffusion formulation, only treats the U-Net as a deterministic backbone and still falls short in performance. In contrast, Lotus systematically analyzes the standard stochastic diffusion formulation for dense prediction and proposes innovations such as the detail preserver to improve accuracy especially in detailed area, finally delivering much better results (Tab. 1). Additionally, Lotus is a stochastic model. In contrast to GenPercept’s deterministic nature, Lotus enables uncertainty predictions. StableNormal (Ye et al., 2024) predicts normal maps through a two-stage process. While the first stage produces coarse normal maps with single-step diffusion, the second stage performs refinement still with iterative diffusion which is computation-intensive. In comparison, Lotus not only achieves fine-grained predictions thanks to our novel detail preserver without extra stages or parameters, but also delivers much superior results (Tab. 2) thanks to our designed diffusion formulation that better fits the pre-trained diffusion for dense prediction.
@@ 3 Preliminaries
Diffusion Formulation for Dense Prediction. Following (Ke et al., 2024) and (Fu et al., 2024), we also formulate dense prediction as an image-conditioned annotation generation task based on Stable Diffusion (Rombach et al., 2022), which performs the diffusion process in low-dimensional latent space for computational efficiency. First, there is a pair of auto-encoders $\{\mathcal{E}(\cdot),\mathcal{D}(\cdot)\}$ trained to map between RGB space and latent space, i.e., $\mathcal{E}(\textbf{x})=\mathbf{z^{x}}$, $\mathcal{D}(\mathbf{z^{x}})\approx\textbf{x}$. The auto-encoder also maps between dense annotations and latent space effectively, i.e., $\mathcal{E}(\textbf{y})=\mathbf{z^{y}}$, $\mathcal{D}(\mathbf{z^{y}})\approx\textbf{y}$ (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024). Following (Ho et al., 2020), Stable Diffusion establishes a pair of forward nosing and reversal denoising processes in latent space. In forward process, Gaussian noise is gradually added at levels $t\in\left[1,T\right]$ into sample $\mathbf{z^{y}}$ to obtain the noisy sample $\mathbf{z^{y}_{t}}$:
@@ 4.1 Parameterization Types
The type of parameterization is a vital configuration, it not only determines the loss function discussed in Sec. 3, but also influences the inference process (Eq. 3). During inference, the predicted clean sample $\mathbf{\hat{z}^{y}_{\tau}}$, a key component in Eq. 3, is calculated according to different parameterizations  [^2] The latest parameterization, $v$-prediction, combines $\epsilon$-prediction and $x_{0}$-prediction, producing results that are intermediate between the two. Please see Sec. D of the supplementary materials for more details..
@@ 4.4 Stochastic Nature of Diffusion Model
One major characteristic of generative models is their stochastic nature, which, in image generation, enables the production of diverse outputs. In perception tasks like dense prediction, this stochasticity has the potential to allow the model generating predictions with uncertainty maps. Specifically, for any input image, we can conduct multiple inferences using different initialization noises and aggregate these predictions to calculate its uncertainty map. Thanks to our systematic analysis and tailored fine-tuning protocol, our method effectively reduces excessive flickering (large variance), only allowing for more accurate uncertainty calculations in naturally uncertain areas, such as the sky, object edges, and fine details (e.g. cat whiskers), as shown in Fig. 9.
@@ 5.1 Experimental Settings
Implementation details. We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), with text conditioning disabled. During training, we fix the time-step $t=1000$. For more details, please see Sec. A.1 of the supplementary materials.
@@
Evaluation Datasets and Metrics. ① For zero-shot affine-invariant depth estimation, we evaluate Lotus on NYUv2 (Silberman et al., 2012), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013) and ETH3D (Schops et al., 2017) using absolute mean relative error (AbsRel), and also report $\delta 1$ and $\delta 2$ values. ② For surface normal prediction, we employ NYUv2, ScanNet, iBims-1 (Koch et al., 2018) and Sintel (Butler et al., 2012) datasets, reporting mean angular error (m.) as well as the percentage of pixels with an angular error below $11.25^{\circ}$ and $30^{\circ}$. Please see Sec. A.2 of the supplementary materials for further details on the evaluation datasets and metrics.
@@ 5.2 Quantitative Comparisons
① For depth estimation (Tab. 1), Lotus-G achieves the best performance compared to all generative baselines. Notice that we only require single step denoising process, also significantly boosting the inference speed as shown in Fig. 3. Lotus-D also performs well, only slightly inferior to DepthAnything. However, it is worthy to notice that Lotus is trained on only 0.059M images compared to DepthAnything’s 62.6M images. ② For normal estimation (Tab. 2), both Lotus-G and Lotus-D outperform all other generative and discriminative methods on zero-shot surface normal estimation with significant margins. Please see Sec. E of the supplementary materials for Qualitative Comparisons.
Table 1: Quantitative comparison on zero-shot affine-invariant depth estimation between Lotus and SoTA methods. The upper section lists discriminative methods, the lower lists generative ones. The best and second best performances are highlighted. Lotus-G outperforms all others methods while Lotus-D is only slightly inferior to DepthAnything. Please note that DepthAnything is trained on 62.6M images while Lotus is only trained on 0.059M images. §indicates results revised by ourselves. denotes the method relies on pre-trained Stable Diffusion.
Method Training NYUv2 (Indoor) KITTI (Outdoor) ETH3D (Various) ScanNet (Indoor) Avg.
@@
DiverseDepth 320K 11.7 87.5 - 19.0 70.4 - 22.8 69.4 - 10.9 88.2 - 9.5
MiDaS 2M 11.1 88.5 - 23.6 63.0 - 18.4 75.2 - 12.1 84.6 - 9.5
LeRes 354K 9.0 91.6 - 14.9 78.4 - 17.1 77.7 - 9.1 91.7 - 7.6
Omnidata 12.2M 7.4 94.5 - 14.9 83.5 - 16.6 77.8 - 7.5 93.6 - 6.4
DPT 1.4M 9.8 90.3 - 10.0 90.1 - 7.8 94.6 - 8.2 93.4 - 5.5
HDN 300K 6.9 94.8 - 11.5 86.7 - 12.1 83.3 - 8.0 93.9 - 5.0
${}^{\star^{\S}}$ 74K 5.6 96.0 99.2 13.0 84.2 - 7.0 95.6 98.8 6.2 96.1 99.1 3.8
DepthAnything V2 62.6M 4.5 97.9 99.3 7.4 94.6 98.6 13.1 86.5 - 4.2 97.8 99.3 2.9
Lotus-D (Ours) 59K 5.3 96.7 99.2 9.3 92.8 98.8 6.8 95.3 98.9 6.0 96.3 99.1 2.5
DepthAnything 62.6M 4.3 98.1 99.6 7.6 94.7 99.2 12.7 88.2 - 4.3 98.1 99.6 2.0
${}^{\star^{\S}}$ 280K 5.6 96.3 99.1 14.4 82.0 96.6 6.6 95.8 98.4 6.4 95.0 98.4 3.3
Marigold${}_{\text{(LCM)}}$§⋆ 74K 6.1 95.8 99.0 9.8 91.8 98.7 6.8 95.6 99.0 6.9 94.6 98.6 2.9
Marigold 74K 5.5 96.4 99.1 9.9 91.6 98.7 6.5 95.9 99.0 6.4 95.2 98.8 1.8
Lotus-G (Ours) 59K 5.4 96.6 99.2 11.3 87.7 97.8 6.2 96.1 99.0 6.0 96.0 99.0 1.5
@@
Method Training NYUv2 (Indoor) ScanNet (Indoor) iBims-1 (Indoor) Sintel (Outdoor) Avg.
Data m.$\downarrow$ $11.25^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ m.$\downarrow$ $11.25^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ m.$\downarrow$ $11.25^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ m.$\downarrow$ $11.25^{\circ}$$\uparrow$ $30^{\circ}$$\uparrow$ Rank
OASIS 110K 29.2 23.8 60.7 32.8 15.4 52.6 32.6 23.5 57.4 43.1 7.0 35.7 7.0
Omnidata 12.2M 23.1 45.8 73.6 22.9 47.4 73.2 19.0 62.1 80.1 41.5 11.4 42.0 5.3
EESNU 2.5M 16.2 58.6 83.5 - - - 20.0 58.5 78.2 42.1 11.5 41.2 4.4
GenPercept 74K 18.2 56.3 81.4 17.7 58.3 82.7 18.2 64.0 82.0 37.6 16.2 51.0 3.7
Omnidata V2 12.2M 17.2 55.5 83.0 16.2 60.2 84.7 18.2 63.9 81.1 40.5 14.7 43.5 3.6
DSINE 160K 16.4 59.6 83.5 16.2 61.0 84.4 17.1 67.4 82.3 34.9 21.5 52.7 1.8
Lotus-D (Ours) 59K 16.8 58.2 83.6 15.3 62.9 85.7 17.7 64.9 82.5 34.6 20.5 55.8 1.6
Marigold‡⋆ 74K 20.9 50.5 - 21.3 45.6 - 18.5 64.7 - - - - 4.2
GeoWizard 280K 18.9 50.7 81.5 17.4 53.8 83.5 19.3 63.0 80.3 40.3 12.3 43.5 3.2
StableNormal 250K 18.6 53.5 81.7 17.1 57.4 84.1 18.2 65.0 82.4 36.7 14.1 50.7 2.0
Lotus-G (Ours) 59K 16.9 59.1 83.2 15.3 64.0 85.2 17.5 66.1 82.7 35.2 19.9 54.8 1.0
@@
$+$ Mixture Dataset (Lotus-G) 59K 5.425 96.597 99.156 11.324 87.692 97.780 6.172 96.077 98.980 6.024 96.026 99.730
-- Noise Input (Lotus-D) 59K 5.334 96.729 99.198 9.334 92.813 98.795 6.846 95.290 98.899 5.982 96.287 99.087
5.3 Ablation Study
As shown in Tab. 3, we conduct ablation studies to validate our designs. Starting with “Direct Adaptation”, we incrementally test the effects of different components, such as parameterization types, the single-step diffusion process, and the detail preserver. Initially, we train the model using only the Hypersim dataset to establish a baseline. We then expand the training dataset using a mixture dataset strategy by including Virtual KITTI, aiming to enhance the model’s generalization ability across different domains. The findings from these ablations validate the effectiveness of our proposed adaptation protocol, demonstrating that each design plays a vital role in optimizing the diffusion models for dense prediction tasks.

### v5

@@ Abstract
Leveraging the visual priors of pre-trained text-to-image diffusion models offers a promising solution to enhance zero-shot generalization in dense prediction tasks. However, existing methods often uncritically use the original diffusion formulation, which may not be optimal due to the fundamental differences between dense prediction and image generation. In this paper, we provide a systemic analysis of the diffusion formulation for the dense prediction, focusing on both quality and efficiency. And we find that the original parameterization type for image generation, which learns to predict noise, is harmful for dense prediction; the multi-step noising/denoising diffusion process is also unnecessary and challenging to optimize. Based on these insights, we introduce Lotus, a diffusion-based visual foundation model with a simple yet effective adaptation protocol for dense prediction. Specifically, Lotus is trained to directly predict annotations instead of noise, thereby avoiding harmful variance. We also reformulate the diffusion process into a single-step procedure, simplifying optimization and significantly boosting inference speed. Additionally, we introduce a novel tuning strategy called detail preserver, which achieves more accurate and fine-grained predictions. Without scaling up the training data or model capacity, Lotus achieves promising performance in zero-shot depth and normal estimation across various datasets. It also enhances efficiency, being significantly faster than most existing diffusion-based methods. Lotus’ superior quality and efficiency also enable a wide range of practical applications, such as joint estimation, single/multi-view 3D reconstruction, etc. Project page: lotus3d.github.io.
Figure 1: We present Lotus, a diffusion-based visual foundation model for dense geometry prediction. With minimal training data, Lotus achieves promising performance in zero-shot depth and normal estimation. “Avg. Rank” indicates the average ranking across all metrics, where lower values are better. Bar length represents the amount of training data used.
@@ 1 Introduction
To validate Lotus, we conduct extensive experiments on two primary geometric dense prediction tasks: zero-shot monocular depth and normal estimation. The results demonstrate that Lotus achieves promising, and even superior, performance on these tasks across a wide range of evaluation datasets. Compared to traditional discriminative methods, Lotus delivers remarkable results with only 59K training samples by effectively leveraging the powerful diffusion priors. Among generative approaches, Lotus also outperforms previous methods in both accuracy and efficiency, being significantly faster than methods like Marigold (Ke et al., 2024) (Fig. 3). Beyond these improvements, Lotus seamlessly supports various applications, such as joint estimation, single/multi-view 3D reconstruction, etc.
@@
  • Based on our insights, we introduce Lotus, a diffusion-based visual foundation model for dense prediction with simple yet effective fine-tuning protocol. Lotus achieves promising performance on both zero-shot monocular depth and surface normal estimation. It also enables a wide range of applications.
@@ 2.2 Generative Models for Dense Perception
Currently, a notable trend involves adopting pre-trained generative models, particularly diffusion models, into dense prediction tasks. Marigold (Ke et al., 2024) and GeoWizard (Fu et al., 2024) directly apply the standard diffusion formulation and the pre-trained parameters, without addressing the inherent differences between image generation and dense prediction, leading to constrained performance. Their efficiency is also severely limited by standard iterative denoising processes and ensemble inferences. In this paper, we propose a novel diffusion formulation tailored to the of dense prediction. Aiming to fully leveraging the pre-trained diffusion’s powerful visual priors, Lotus enables more accurate and efficient predictions, finally achieving promising performance.
More recent works, GenPercept (Xu et al., 2024) and StableNormal (Ye et al., 2024), also adopted single-step diffusion. However, GenPercept (Xu et al., 2024) first removes noise input for deterministic characteristic based on DMP (Lee et al., 2024), and then adopts one-step strategy to avoid surface texture interference. It lacks systematic analysis of the diffusion formulation, only treats the U-Net as a deterministic backbone and still falls short in performance. In contrast, Lotus systematically analyzes the standard stochastic diffusion formulation for dense prediction and proposes innovations such as the detail preserver to improve accuracy especially in detailed area, finally delivering much better results (Tab. 1). Additionally, Lotus is a stochastic model. In contrast to GenPercept’s deterministic nature, Lotus enables uncertainty predictions. StableNormal (Ye et al., 2024) predicts normal maps through a two-stage process. While the first stage produces coarse normal maps with single-step diffusion, the second stage performs refinement still with iterative diffusion which is computation-intensive. In comparison, Lotus not only achieves fine-grained predictions thanks to our novel detail preserver without extra stages or parameters, but also delivers much superior results (Tab. 2) thanks to our designed diffusion formulation that better fits the pre-trained diffusion for dense prediction. Recently, a concurrent work, Diffusion-E2E-FT (Garcia et al., 2024), has also achieved promising results in a single step. Its main contribution lies in addressing the issue where Marigold (Ke et al., 2024) and similar models (Fu et al., 2024) use inconsistent pairings of time-step and noise, resulting in poor predictions. By setting the “time-step spacing” to “trailing” mode in schedulers, it prevents “GT” signal leakage during inference, improving accuracy. While the performance of Lotus-D and Diffusion-E2E-FT is similar, Lotus is based on a systematic analysis of stochastic diffusion for dense prediction, with innovations like the detail preserver to enhance accuracy, particularly in detailed areas. Additionally, unlike the deterministic Diffusion-E2E-FT, Lotus (Lotus-G) is a stochastic model that enables uncertainty predictions.
@@ 3 Preliminaries
Diffusion Formulation for Dense Prediction. Following (Ke et al., 2024) and (Fu et al., 2024), we also formulate dense prediction as an image-conditioned annotation generation task based on Stable Diffusion (Rombach et al., 2022), which performs the diffusion process in low-dimensional latent space for computational efficiency. First, the auto-encoder, which consists an encoder $\mathcal{E}(\cdot)$ and a decoder $\mathcal{D}(\cdot)$, is trained to map between RGB space and latent space, i.e., $\mathcal{E}(\textbf{x})=\mathbf{z^{x}}$, $\mathcal{D}(\mathbf{z^{x}})\approx\textbf{x}$. The auto-encoder also maps between dense annotations and latent space effectively, i.e., $\mathcal{E}(\textbf{y})=\mathbf{z^{y}}$, $\mathcal{D}(\mathbf{z^{y}})\approx\textbf{y}$ (Ke et al., 2024; Fu et al., 2024; Xu et al., 2024; Ye et al., 2024). Following (Ho et al., 2020), Stable Diffusion establishes a pair of forward nosing and reversal denoising processes in latent space. In forward process, Gaussian noise is gradually added at levels $t\in\left[1,T\right]$ into sample $\mathbf{z^{y}}$ to obtain the noisy sample $\mathbf{z^{y}_{t}}$:
@@ 4.1 Parameterization Types
The type of parameterization is crucial, it not only determines the loss function discussed in Sec. 3, but also influences the inference process (Eq. 3). During inference, the predicted clean sample $\mathbf{\hat{z}^{y}_{\tau}}$, a key component in Eq. 3, is calculated according to different parameterizations  [^2] The latest parameterization, $v$-prediction, combines $\epsilon$-prediction and $x_{0}$-prediction, producing results that are intermediate between the two. Please see Sec. D of the supplementary materials for more details..
@@ 4.4 Stochastic Nature of Diffusion Model
One major characteristic of generative models is their stochastic nature, which, in image generation, enables the production of diverse outputs. In perception tasks like dense prediction, this stochasticity has the potential to allow the model generating predictions with uncertainty maps. Specifically, for any input image, we can conduct multiple inferences using different initialization noises and aggregate these predictions to calculate its uncertainty map. Thanks to our systematic analysis and tailored fine-tuning protocol, our method effectively
reduces excessive flickering (large variance), only allowing for more accurate uncertainty calculations in naturally uncertain areas, such as the sky, object edges, and fine details (e.g. cat whiskers), as shown in Fig. 9.
@@ 5.1 Experimental Settings
Implementation details. We implement Lotus based on Stable Diffusion V2 (Rombach et al., 2022), without text conditioning. During training, we fix the time-step $t=1000$. For depth estimation, we predict in disparity space, i.e., $d=1/d^{\prime}$, where $d$ represents the values in disparity space and $d^{\prime}$ denotes the true depth. For more details, please see Sec. A.1 of the supplementary materials.
@@
Evaluation Datasets and Metrics. ① For zero-shot affine-invariant depth estimation, we evaluate Lotus on NYUv2 (Silberman et al., 2012), ScanNet (Dai et al., 2017), KITTI (Geiger et al., 2013), ETH3D (Schops et al., 2017), and DIODE (Vasiljevic et al., 2019) using absolute mean relative error (AbsRel), and also report $\delta 1$ and $\delta 2$ values. ② For surface normal prediction, we employ NYUv2, ScanNet, iBims-1 (Koch et al., 2018), Sintel (Butler et al., 2012), and OASIS (Chen et al., 2020) datasets, reporting mean angular error (m.) as well as the percentage of pixels with an angular error below $11.25^{\circ}$ and $30^{\circ}$. Please see Sec. A.2 of the supplementary materials for further details on the evaluation datasets and metrics.
@@ 5.2 Quantitative Comparisons
① For depth estimation (Tab. 1), Lotus-G demonstrates promising performance across all evaluation datasets, achieving the overall best rank compared to other generative baselines. Notice that we only require single step denoising process, significantly boosting the inference speed as shown in Fig. 3. Lotus-D also performs well, achieving comparable results to DepthAnything series. It is worthy to notice that Lotus is trained on only 0.059M images compared to DepthAnything’s 62.6M images. ② For normal estimation (Tab. 2), both Lotus-G and Lotus-D outperform all other generative and discriminative methods on zero-shot surface normal estimation with significant margins. Please see Sec. H of the supplementary materials for Qualitative Comparisons.
5.3 Ablation Study
As shown in Tab. 3, we conduct ablation studies to validate our designs. Starting with “Direct Adaptation”, we incrementally test the effects of different components, such as parameterization types, the single-step diffusion process, and the detail preserver. Initially, we train the model using only the Hypersim dataset to establish a baseline. We then expand the training dataset using a mixture dataset strategy by including Virtual KITTI, aiming to enhance the model’s generalization ability across different domains. For depth estimation, we further train the model in the disparity space to improve the accuracy. The findings from these ablations validate the effectiveness of our proposed adaptation protocol, demonstrating that each design plays a vital role in optimizing the diffusion models for dense prediction tasks.
Table 1: Quantitative comparison on zero-shot affine-invariant depth estimation between Lotus and SoTA methods. The upper section lists discriminative methods, the lower lists generative ones. The best and second best performances are highlighted. Lotus-G outperforms all others methods while Lotus-D is only slightly inferior to DepthAnything. §indicates results revised by ourselves, following Marigold (Ke et al., 2024). denotes the method relies on pre-trained Stable Diffusion.
Method Training NYUv2 (Indoor) KITTI (Outdoor) ETH3D (Various) ScanNet (Indoor) DIODE (Various) Avg
@@ 5.3 Ablation Study
DiverseDepth 320K 11.7 87.5 - 19.0 70.4 - 22.8 69.4 - 10.9 88.2 - 37.6 63.1 - 10.6
MiDaS 2M 11.1 88.5 - 23.6 63.0 - 18.4 75.2 - 12.1 84.6 - 33.2 71.5 - 10.2
LeRes 354K 9.0 91.6 - 14.9 78.4 - 17.1 77.7 - 9.1 91.7 - 27.1 76.6 - 7.8
Omnidata 12.2M 7.4 94.5 - 14.9 83.5 - 16.6 77.8 - 7.5 93.6 - 33.9 74.2 - 7.5
DPT 1.4M 9.8 90.3 - 10.0 90.1 - 7.8 94.6 - 8.2 93.4 - 18.2 75.8 - 5.8
HDN 300K 6.9 94.8 - 11.5 86.7 - 12.1 83.3 - 8.0 93.9 - 24.6 78.0 - 5.3
${}^{\star^{\S}}$ 74K 5.6 96.0 99.2 13.0 84.2 97.2 7.0 95.6 98.8 6.2 96.1 99.1 35.7 75.6 86.6 4.9
Diffusion-E2E-FT 74K 5.4 96.5 99.1 9.6 92.1 98.0 6.4 95.9 98.7 5.8 96.5 98.8 30.3 77.6 87.9 3.6
DepthAnything V2 62.6M 4.5 97.9 99.3 7.4 94.6 98.6 13.1 86.5 97.5 4.2 97.8 99.3 26.5 73.4 87.1 3.5
Lotus-D (Ours) 59K 5.1 97.2 99.2 8.1 93.1 98.7 6.1 97.0 99.1 5.5 96.5 99.0 22.8 73.8 86.2 3.0
DepthAnything 62.6M 4.3 98.1 99.6 7.6 94.7 99.2 12.7 88.2 98.3 4.3 98.1 99.6 26.0 75.9 87.5 2.4
${}^{\star^{\S}}$ 280K 5.6 96.3 99.1 14.4 82.0 96.6 6.6 95.8 98.4 6.4 95.0 98.4 33.5 72.3 86.5 3.3
Marigold${}_{\text{(LCM)}}$⋆§ 74K 6.1 95.8 99.0 9.8 91.8 98.7 6.8 95.6 99.0 6.9 94.6 98.6 30.7 77.5 89.3 2.9
Marigold 74K 5.5 96.4 99.1 9.9 91.6 98.7 6.5 95.9 99.0 6.4 95.2 98.8 30.8 77.3 88.7 2.1
Lotus-G (Ours) 59K 5.4 96.8 99.2 8.5 92.2 98.4 5.9 97.0 99.2 5.9 95.7 98.8 22.9 72.9 86.0 1.3
@@
Method Training NYUv2 (Indoor) ScanNet (Indoor) iBims-1 (Indoor) Sintel (Outdoor) OASIS (Various) Avg.
[…]
OASIS 110K 29.2 23.8 60.7 32.8 15.4 52.6 32.6 23.5 57.4 43.1 7.0 35.7 - - - 7.8
Omnidata 12.2M 23.1 45.8 73.6 22.9 47.4 73.2 19.0 62.1 80.1 41.5 11.4 42.0 24.9 31.0 71.4 5.9
EESNU 2.5M 16.2 58.6 83.5 - - - 20.0 58.5 78.2 42.1 11.5 41.2 27.7 24.0 66.6 5.8
GenPercept§⋆ 74K 18.2 56.3 81.4 17.7 58.3 82.7 18.2 64.0 82.0 37.6 16.2 51.0 26.3 26.9 71.1 4.9
Omnidata V2 12.2M 17.2 55.5 83.0 16.2 60.2 84.7 18.2 63.9 81.1 40.5 14.7 43.5 24.2 27.7 74.2 4.4
DSINE 160K 16.4 59.6 83.5 16.2 61.0 84.4 17.1 67.4 82.3 34.9 21.5 52.7 24.4 28.8 72.0 3.1
Diffusion-E2E-FT§⋆ 74K 16.5 60.4 83.1 14.7 66.1 85.1 16.1 69.7 83.9 33.5 22.3 53.5 23.2 29.4 74.5 1.9
Lotus-D (Ours) 59K 16.2 59.8 83.9 14.7 64.0 86.1 17.1 66.4 83.0 32.3 22.4 57.0 22.3 31.8 76.1 1.4
Marigold‡⋆ 74K 20.9 50.5 - 21.3 45.6 - 18.5 64.7 - - - - - - - 3.6
GeoWizard§⋆ 280K 18.9 50.7 81.5 17.4 53.8 83.5 19.3 63.0 80.3 40.3 12.3 43.5 25.2 23.4 68.1 3.1
StableNormal§⋆ 250K 18.6 53.5 81.7 17.1 57.4 84.1 18.2 65.0 82.4 36.7 14.1 50.7 26.5 23.5 68.7 2.1
Lotus-G (Ours) 59K 16.5 59.4 83.5 15.1 63.9 85.3 17.2 66.2 82.7 33.6 21.0 53.8 22.7 29.4 75.8 1.0
@@
$+$ Mixture Dataset 59K 5.425 96.597 99.156 11.324 87.692 97.780 6.172 96.077 98.980 6.024 96.026 99.730
$\hookrightarrow$ -- Noise Input 59K 5.334 96.729 99.198 9.334 92.813 98.795 6.846 95.290 98.899 5.982 96.287 99.087
$+$ Disparity Space (Lotus-G) 59K 5.379 96.736 99.155 8.521 92.206 98.374 5.878 97.024 99.233 5.925 95.727 98.839
  $\hookrightarrow$ -- Noise Input (Lotus-D) 59K 5.123 97.182 99.134 8.117 93.097 98.654 6.147 96.964 99.077 5.494 96.534 99.039